OpenStudy (anonymous):

find the equation of a straight line with slope -2 and intersecting the x-axis at a distance of 3 units to the left of the origin

OpenStudy (gorv):



OpenStudy (gorv):

y=mx+c m=slope

OpenStudy (gorv):

y=-2x+c

OpenStudy (gorv):

3 unit to left of origin means x=-3

OpenStudy (gorv):

at x=-3 y=0

OpenStudy (gorv):

0=-2*(-3)+c

OpenStudy (gorv):

0=6+c c=-6

OpenStudy (gorv):

y=-2x-6

OpenStudy (anonymous):

Oh! Thank you.

OpenStudy (gorv):

welcome :)